# Tight MIP Formulations for Optimal Operation and Investment of Storage Including Reserves

### v2

@@ V Conclusions
The presented case studies illustrate the improved ability of the LP relaxations of the tighter MIP formulations to prevent simultaneous charging and discharging in a multi-period problem, compared to other LP models. Furthermore, we illustrated that the tighter MIP formulation can indeed positively affect the solving time of the model, also for multi-period problems.
@@
We use the method described in III-A to obtain the convex hull. First, we write two disjunctive sets of constraints in Section -A1. We then obtain the convex hull of these two disjunctive sets of constraints in Section -A2, as explained by Balas [25]. Lastly, we obtain the convex hull in the dimension of the original formulation by eliminating the extra variable using the Fourier-Motzkin elimination procedure in Section -A3.
@@
[…]
@@
We can then obtain the convex hull of these two disjunctive sets of constraints, as described Balas [25] in the following way. We rename all variables from one set to $\square^{1}$ and variables from the other to $\square^{2}$, and we multiply all parameters in one set with $\delta^{1}$, and the others with $\delta^{2}$. We also need to include the additional constraint $\delta^{1}+\delta^{2}=1$. This gives us the constraints in (24), which describe the convex hull of the solutions to the BO-MIP model for one time period.
@@
$\underline{E}\delta\leq e$ $-\frac{1}{\eta^{\mathrm{D}}}p^{\mathrm{D}}\Delta-\underline{E}(1-\delta)$
[…]
@@
As explained in Section IV, we used the unit commitment and transmission expansion planning case studies by Arroyo et al. [6], and adapted them to include reserves and investment decisions. For a full explanation of the case studies, we refer to the original paper. In this section, we explain how we adapted the case studies.
An overview of all the parameter values in the original UC case study, as well as in our adapted case studies (used to obtain the results in Tables I and II can be found in Table V. For the regular optimal operation problem, we slightly adjusted the charging/discharging capacities. As explained at the start of Section III, the LP relaxations of our formulations describe the convex hull of the solutions to the MIP models for one time period under the assumption that $\overline{P^{\mathrm{C}}}\leq\frac{1}{\eta^{\mathrm{C}}\Delta}(\overline{E}-\underline{E})$ and $\overline{P^{\mathrm{D}}}\leq\frac{\eta^{\mathrm{D}}}{\Delta}(\overline{E}-\underline{E})$. In the original case study, the values of $\overline{P^{\mathrm{C}}}$ and $\overline{P^{\mathrm{D}}}$ do not satisfy this, so we changed these values, as can be seen in Table VI. Note that this does not affect the optimal mixed-integer solution.
@@
$R^{+}$ $\frac{1}{\eta^{\mathrm{C}}\Delta}(\overline{E}-\underline{E})$
$R^{-}$ $\frac{\eta^{\mathrm{D}}}{\Delta}(\overline{E}-\underline{E})$
@@
Constraints (27a) and (27b) together with (2c) and (2d) can be rewritten as (28b) and (28c). This results in the BOF-MIP model (Basic Operation with Flexible Reserves MIP model), as presented by Momber et al.[19]. Now, (28b) and (28c) ensure that the down reserves $r^{-}_{t}$ and up reserves $r^{+}_{t}$, respectively, are essentially bounded by $\overline{P^{\mathrm{D}}}+\overline{P^{\mathrm{C}}}$. Thus, this formulation does fully exploit the flexibility of the storage unit, as illustrated in Figure 1. Lastly, the capacity of the reserves is also bounded by the current state of charge, as modeled by (28a). We provide a tighter formulation of this model in the online companion[20, Section 4].

### v3

@@ V Conclusions
The presented case studies illustrate the improved ability of the LP relaxations of the tighter MIP formulations to prevent simultaneous charging and discharging in a multi-period problem, compared to other LP models. Furthermore, we illustrated that the tighter MIP formulation could indeed positively affect the solving time of the model, also for multi-period problems, though this is case study dependent.
@@
We use the method described in III-A to obtain the convex hull. First, we write two disjunctive sets of constraints in Section -A1. We then obtain the convex hull of these two disjunctive sets of constraints in Section -A2, as explained by Balas [balasDisjunctiveProgrammingHierarchy1985]. Lastly, we obtain the convex hull in the dimension of the original formulation by eliminating the extra variable using the Fourier-Motzkin elimination procedure in Section -A3.
@@
$\begin{array}[]{r @{} l}\text{Charging:}\\ \delta&\ =1\\ p^{\mathrm{D}}&\ =0\\ \cancel{\underline{E}\leq}\ e+\eta^{\mathrm{C}}p^{\mathrm{C}}\Delta&\ \leq\overline{E}\\ \underline{E}\leq e&\ \cancel{\leq\overline{E}}\\ p^{\mathrm{C}}&\ \leq\overline{P^{\mathrm{C}}}\\ e&\ \cancel{\geq 0}\\ p^{\mathrm{C}}&\ \geq 0\end{array}$ $\quad\mathrm{or}\quad\begin{array}[]{r@{}l}\text{Discharging:}\\ \delta&\ =0\\ p^{\mathrm{C}}&\ =0\\ \underline{E}\leq e-\frac{1}{\eta^{\mathrm{D}}}p^{\mathrm{D}}\Delta&\ \cancel{\leq\overline{E}}\\ \cancel{\underline{E}\leq}\ e&\ \leq\overline{E}\\ p^{\mathrm{D}}&\ \leq\overline{P^{\mathrm{D}}}\\ e&\ \cancel{\geq 0}\\ p^{\mathrm{D}}&\ \geq 0\end{array}$ (23)
@@
We can then obtain the convex hull of these two disjunctive sets of constraints, as described Balas [balasDisjunctiveProgrammingHierarchy1985] in the following way. We rename all variables from one set to $\square^{1}$ and variables from the other to $\square^{2}$, and we multiply all parameters in one set with $\delta^{1}$, and the others with $\delta^{2}$. We also need to include the additional constraint $\delta^{1}+\delta^{2}=1$. This gives us the constraints in (24), which describe the convex hull of the solutions to the BO-MIP model for one time period.
@@
$\underline{E}\delta\leq e\$$-\frac{1}{\eta^{\mathrm{D}}}p^{\mathrm{D}}\Delta-\underline{E}(1-\delta)$
$\Rightarrow e\geq\underline{E}+\frac{1}{\eta^{\mathrm{D}}}p^{\mathrm{D}}\Delta$ $\text{ in CH}.$ (26b)
@@
As explained in Section IV, we used the unit commitment and transmission expansion planning case studies by Arroyo et al. [arroyoUseConvexModel2020], and adapted them to include reserves and investment decisions. For a full explanation of the case studies, we refer to the original paper. In this section, we explain how we adapted the case studies.
An overview of all the parameter values in the original UC case study, as well as in our adapted case studies (used to obtain the results in Tables I and II can be found in Table V. For the regular optimal operation problem, we slightly adjusted the charging/discharging capacities. As explained at the end of Section III-A, the LP relaxations of our formulations describe the convex hull of the solutions to the MIP models for one time period under the assumption that $\overline{P^{\mathrm{C}}}\leq\frac{1}{\eta^{\mathrm{C}}\Delta}(\overline{E}-\underline{E})$ and $\overline{P^{\mathrm{D}}}\leq\frac{\eta^{\mathrm{D}}}{\Delta}(\overline{E}-\underline{E})$. In the original case study, the values of $\overline{P^{\mathrm{C}}}$ and $\overline{P^{\mathrm{D}}}$ do not satisfy this, so we changed these values, as can be seen in Table VI. Note that this does not affect the optimal mixed-integer solution.
@@
$R^{+}$ - - $\frac{1}{\eta^{\mathrm{C}}\Delta}(\overline{E}-\underline{E})$
$R^{-}$ - - $\frac{\eta^{\mathrm{D}}}{\Delta}(\overline{E}-\underline{E})$
@@
Constraints (27a) and (27b) together with (2c) and (2d) can be rewritten as (28b) and (28c). This results in the BOF-MIP model (Basic Operation with Flexible Reserves MIP model), as presented by Momber et al.[momberPEVStorageMultiBus2014]. Now, (28b) and (28c) ensure that the down reserves $r^{-}_{t}$ and up reserves $r^{+}_{t}$, respectively, are essentially bounded by $\overline{P^{\mathrm{D}}}+\overline{P^{\mathrm{C}}}$. Thus, this formulation does fully exploit the flexibility of the storage unit, as illustrated in Figure 1. Lastly, the capacity of the reserves is also bounded by the current state of charge, as modeled by (28a). We provide a tighter formulation of this model in the online companion[elgersmaOnlineCompanionTight2024, Section 4].